OpenStudy (anonymous):

PLEASE HELP!!! Suppose you make 90% of your free throws and you attempt 3 free throws. Use the bionomial theorem to calculate each probability. a.) you don't make any of them b.) you make 1 c.) you make 2 d.) you make all 3

OpenStudy (amistre64):

whats the thrm tell us?

OpenStudy (anonymous):

P(x)= vnC vx P^xq^n-x

OpenStudy (anonymous):

but I don't understand how to find C?

OpenStudy (amistre64):

C is a notation for combinatorics

OpenStudy (amistre64):

n C r means:\[\frac{n!}{r!(n-r)!}\]

OpenStudy (amistre64):

another way of look at it is: (p+q)^n ; such that p+q = 1 \[(.90 + .10)^3 = 1(.90)^3(.10)^0+ 3(.90)^2(.10)^1 + 3(.90)^1(.10)^2+ 1(.90)^0(.10)^3\] its a binomial cumulative distribution function

OpenStudy (anonymous):

thank you so much! can you just show me what the equation would look like for A? I should be able to solve it i'm just having trouble writing the equation out.

OpenStudy (amistre64):

\[ \underbrace{1(.90)^3(.10)^0}_{P(3)}+ \underbrace{3(.90)^2(.10)^1}_{P(2)} + \underbrace {3(.90)^1(.10)^2}_{P(1)}+ \underbrace{1(.90)^0(.10)^3}_{P(0)}\]

OpenStudy (amistre64):

the nCr is the coefficients of a pascal triangle, ^3 = 1 3 3 1 \[\binom{3}{0}=3C0=1\] \[\binom{3}{1}=3C1=3\] \[\binom{3}{2}=3C2=3\] \[\binom{3}{3}=3C3=1\]

OpenStudy (amistre64):

P(0) is the probability that we make none of them ....

OpenStudy (anonymous):

thank you SOO much!!

OpenStudy (amistre64):

youre welcome